Monty Corporation operates a retail computer store. To improve delivery services to customers, the company purchases four new trucks on April 1,2017. The terms of acquisition for each truck are described below. 1. 2. Truck #1 has a list price of $ 18,750 and is acquired for a cash payment of $ 17,375. Truck #2 has a list price of $ 20,000 and is acquired for a down payment of $ 2,500 cash and a zero-interest-bearing note with a face amount of $ 17,500. The note is due April 1, 2018, Monty would normally have to pay interest at a rate of 9% for such a borrowing, and the dealership has an incremental borrowing rate of 8%. Truck #3hasa ist price of $20 inventory system Truck #4 has a list price of $ 17 500. It is acquired in exchange for 940 shares of common stock in Monty Corporation. The stock has a par value per share of $ 10 and a market price of $ 13 per share. 3. .lt s acquired n exchange fraco puter syster that Monty cares in inventoryThe computer syste cost OO andsnorma ysodbyMonty or 19. OLMontus sa petual 4. Prepare the appropriate ournal entries for the above transactions for Monty Corporation Round present value actors to 5 decimal places, eふ0 5258 7 and final answers to 0 decimal places, e account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) 5,275. Credit No. Account Titles and Explanation 1. Debit Credit trucks 17375 cash 17375 2. 18750 Discount on Notes Payable 2695 cash 2500 Notes Payable 17500 3. 4.
Answer #1
No. Account titles and explanation Debit Credit
1 Trucks 17,375
Cash 17,375
2 Trucks 18,555
Discount on Notes Payable 1,445
Cash 2,500
Notes payable 17,500
3 Trucks 19,000
Cost of goods sold 15,000
Inventory 15,000
Sales Revenue 19,000
4 Trucks 12,220
Common stock 9,400
Paid in capital in excess of par common stock 2,820

Calculations and explanations:

1. Transaction 1 is entered at the cash payment amount.

2. For transaction 2 the value of trucks = 17500*1/1.09 = 17500*0.91743 = 16,055.03. Thus value of trucks = 16055.03+2500 = 18555

Discount on notes payable = 2500+17500-16055.03 = 1444.97 or 1445

4. 12220 = 940*13

9400 = 940*10 and 2820 = 940*(13-10)